# Covariance scanning for adaptively optimal change point detection in high-dimensional linear models

### v2

@@ 3.2 Dense regime
The construction of $T_{k}$ is motivated by the estimator of the explained variance proposed by Dicker, (2014) and Verzelen and Gassiat, (2018) and indeed, $T_{\theta}$ approximates $\tfrac{\theta(n-\theta)}{n}|\bm{\Sigma}\bm{\delta}|_{2}^{2}$, which follows from the following general result.
@@
In contrast to $\bar{T}_{k}$ in (3), the statistic $T_{k}$ involves the $\ell_{2}$-aggregation of the contrast between the local sample covariance before and after $k$, and thus lends itself naturally to the detection of the change point in the dense regime. We can write $T_{k}=\sum_{i\in[n]}(\gamma_{i}-\bar{\gamma})\bigl{(}\mathbf{w}_{i}^{\top}\widetilde{\mathbf{Y}}(k)\bigr{)}^{2}$, where $(\gamma_{i},\mathbf{w}_{i}),\,i\in[n]$, denote the pairs of eigenvalues and eigenvectors of $\mathbf{X}\mathbf{X}^{\top}$ and $\bar{\gamma}=n^{-1}\sum_{i\in[n]}\gamma_{i}$. This representation reveals the importance of ‘centring’ of $\mathbf{X}\mathbf{X}^{\top}$ in (6), which enables $T_{k}$ to amplify when $\mathsf{E}(\widetilde{\mathbf{Y}}(k)|\mathbf{X})$ is more aligned with $\mathbf{w}_{i}$’s associated with larger $\gamma_{i}$’s.
@@
Under the AMOC model (1), let $p\geq n$ and $\varpi_{\text{\tiny Q}}\asymp\log\log^{3}(n)$. Suppose either that there is a change point, with $q=1$, $\Delta\geq 2\varpi_{\text{\tiny Q}}$ and
@@
$\mathsf{P}\left\{\widehat{q}_{\text{\tiny Q}}=q\,;\,|\widehat{\theta}_{\text{\tiny Q}}-\theta|\leq c_{1}|\bm{\Sigma}\bm{\delta}|_{2}^{-2}\|\bm{\Sigma}\|\Psi^{2}\sqrt{p\log\log(n)}\right\}\geq 1-c_{2}\log^{-1}(n),$
[…]
With $\sigma_{*}$ denoting the smallest non-zero eigenvalue of $\bm{\Sigma}$, it holds that
$|\bm{\Sigma}\bm{\delta}|_{2}^{2}\geq\sigma_{*}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}.$
Thus, QcScan is minimax near-optimal up to a multiplicative factor of $\sqrt{\log\log(n)}$, in the dense regime where $\mathfrak{s}\geq\sqrt{p}$ (cf. Table 1), provided that $\sigma_{*}$ is bounded away from zero (see also the discussions below Theorem 3). When $\bm{\Sigma}$ is known, pre-rotating $\mathbf{x}_{t}$ allows us to re-write (8) without the dependence on $\sigma_{*}$, as
@@
    As in Verzelen and Gassiat, (2018), we focus on the high-dimensional setting by assuming that $p\geq n$ in Proposition 4. In the low-dimensional setting with $p<n$, the conclusion of the proposition becomes $\mathsf{P}\{|T_{k}-f_{k}|\geq C_{0}\|\bm{\Sigma}\|\Psi^{2}\sqrt{nz}\}\leq C_{1}e^{-z}$ and accordingly, the results in Theorem 5 no longer depend on $p$.
@@
    The proof of Proposition 4 follows from the exponential concentration inequalities for Gaussian chaos (Adamczak and Wolff,, 2015). We may relax the Gaussianity assumption to permit sub-Weibull distributions by means of Theorems 1.3 and 1.5 of Götze et al., (2021), and thus extend the guarantees of QcScan to non-Gaussian settings.
@@ 3.3.1 When the differential parameter is sparse
Directly imposing sparsity on $\bm{\delta}$, Cho et al., (2025) proposed $\ell_{1}$-penalised estimator for its estimation, which is obtained as
@@
Suppose that either (i) $\widehat{\bm{\delta}}$ in (9) is obtained with $\widehat{\theta}=\widehat{\theta}_{\text{\tiny M}}$ and the conditions in Theorem 3 are met, or (ii) $\widehat{\bm{\delta}}$ is obtained with $\widehat{\theta}=\widehat{\theta}_{\text{\tiny Q}}$ and the conditions in Theorem 5 are met. Let $\underline{\sigma}$ denote the smallest eigenvalue of $\bm{\Sigma}$. Provided that $n\geq 4C_{\text{RE}}\underline{\sigma}^{-2}\log(p\vee n)$ and $\lambda=C_{\lambda}\|\bm{\Sigma}\|^{1/2}\Psi\sqrt{\log(p\vee n)}$ for some constants $C_{\text{RE}},C_{\lambda}\in(0,\infty)$, we have under (i),
$\bigl{|}\widehat{\bm{\delta}}-\bm{\delta}\bigr{|}_{2}$ $\lesssim\frac{\|\bm{\Sigma}\|^{1/2}\Psi\sqrt{\mathfrak{s}_{\delta}\log(p\vee n)}}{\underline{\sigma}\sqrt{\Delta}}\text{ \ and \ }\bigl{|}\widehat{\bm{\delta}}-\bm{\delta}\bigr{|}_{1}\lesssim\frac{\|\bm{\Sigma}\|^{1/2}\Psi\mathfrak{s}_{\delta}\sqrt{\log(p\vee n)}}{\underline{\sigma}\sqrt{\Delta}},$
@@
$\bigl{|}\widehat{\bm{\delta}}-\bm{\delta}\bigr{|}_{2}$ $\lesssim\frac{\|\bm{\Sigma}\|^{1/2}\Psi\max\{\sqrt{\mathfrak{s}_{\delta}\log(p\vee n)},\bigl{(}p\log\log(n)\bigr{)}^{1/4}\}}{\underline{\sigma}\sqrt{\Delta}}\text{ \ and \ }$
$\bigl{|}\widehat{\bm{\delta}}-\bm{\delta}\bigr{|}_{1}$ $\lesssim\frac{\|\bm{\Sigma}\|^{1/2}\Psi\sqrt{\mathfrak{s}_{\delta}}\max\{\sqrt{\mathfrak{s}_{\delta}\log(p\vee n)},\bigl{(}p\log\log(n)\bigr{)}^{1/4}\}}{\underline{\sigma}\sqrt{\Delta}},$
with probability at least $1-c_{2}\log^{-1}(n)-c^{\prime}(p\vee n)^{-1}$, for some universal constant $c^{\prime}\in(0,\infty)$ and $\bar{c}_{2}$ and $c_{2}$ in Theorems 3 and 5.
For the consistency of $\widehat{\bm{\delta}}$, we make the stronger assumption requiring $\bar{\sigma}>0$, compared to that required for the consistency in change point detection involving $\sigma_{*}$, the smallest non-zero eigenvalue of $\bm{\Sigma}$. In the inherently sparse regime with $\widehat{\theta}=\widehat{\theta}_{\text{\tiny M}}$, the consistency of $\widehat{\bm{\delta}}$ is analogous to that of the Lasso estimator of the regression parameter (see e.g. Bickel et al.,, 2009), with $\Delta=\min(\theta,n-\theta)$ taking the role of sample size. Cho et al., (2025) additionally propose a bias-corrected estimator of $\bm{\delta}$ and establish a non-asymptotic Gaussian approximation result, which enables simultaneous inference about all $p$ coordinates of $\bm{\delta}$. Later in Section 4.2, we devise a change point estimator which, based on scanning the projected covariance as $\widehat{\bm{\delta}}^{\top}\mathbf{X}^{\top}\widetilde{\mathbf{Y}}(k)$, attains a refined rate of estimation that is inversely proportional to $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$.
@@ 3.3.2 For the differential parameter of arbitrary sparsity
Generally, it is infeasible to infer $\bm{\delta}$, or even the regression parameter under stationarity without any assumption on its sparsity. Yet it is still of interest to learn the strength of the change measured by $V_{\circ}:=\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$, a quantity closely related to the signal-to-noise ratio or the explained variance in linear models. In particular, $\sigma^{-2}V_{\circ}$ determines the total variation distance between the joint distributions of $(\mathbf{x}_{t},Y_{t})$ with and without a change, see Lemma 2. Motivated by Proposition 4 and Theorem 5, we propose the following estimator
@@
with probability at least $1-c^{\prime\prime}\log^{-1}(n)$, with $c^{\prime\prime}\in(0,\infty)$ depending only on $C_{1}$ in Proposition 4.
The RHS of (10) tends to zero as the gap between the two sides of (8) increases. If $\bm{\Sigma}$ is known, we can pre-rotate the data so that $\widehat{V}$ estimates $V_{\circ}$ directly. In the absence of such knowledge, we conjecture that it is not possible to estimate $V_{\circ}$ without imposing any constraint on the sparsity of $\bm{\delta}$, which is supported by the impossibility result derived in Theorem 4.5 of Verzelen and Gassiat, (2018) for the related problem of estimating the explained variance in stationary linear models.
@@ 4.1 Adaptively optimal change point detection
Section 3 demonstrates that McScan and QcScan are minimax (near-)optimal in sparse and dense regimes, respectively. Each regime is determined by whether $\mathfrak{s}<\sqrt{p}$ or not, where $\mathfrak{s}$ is typically unknown in practice. This calls for a method that adapts to unknown sparsity levels. We propose to combine the outputs from McScan and QcScan, which leads to a procedure that detects a change whenever either of the detection lower bounds in (5) and (8) is met. Further, the resulting change point estimator inherits the localisation property from either $\widehat{\theta}_{\text{\tiny M}}$ or $\widehat{\theta}_{\text{\tiny Q}}$ that attains the better rate of estimation. In short, referred to as OcScan (optimal covariance scanning), the combined method adaptively enjoys the optimality of McScan and QcScan without the explicit knowledge of $\mathfrak{s}$.
It is natural to determine the existence of the change point by $\widehat{q}=\max\{\mathbb{I}_{\{\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}>\zeta_{\text{\tiny M}}\}},\,\mathbb{I}_{\{T_{\widehat{\theta}_{\text{\tiny Q}}}>\zeta_{\text{\tiny Q}}\}}\}$, which ensures that a change is detected regardless of the regime determined by $\mathfrak{s}$. If $\widehat{q}=0$, we set ${\widehat{\theta}_{\text{\tiny O}}}=n$, while if $\widehat{q}=1$ (with the convention that $\bar{T}_{n}=T_{n}=0$),
$\widehat{\theta}_{\text{\tiny O}}=\begin{cases}\widehat{\theta}_{\text{\tiny M}}&\text{if \ }\mathbb{I}_{\{\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}>\zeta_{\text{\tiny M}}\}}\cdot\mathbb{I}_{\{T_{\widehat{\theta}_{\text{\tiny Q}}}\leq\zeta_{\text{\tiny Q}}\}}=1\text{ \ or \ }C_{\text{\tiny M}/\text{\tiny Q}}>1,\\ \widehat{\theta}_{\text{\tiny Q}}&\text{if \ }\mathbb{I}_{\{\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}\leq\zeta_{\text{\tiny M}}\}}\cdot\mathbb{I}_{\{T_{\widehat{\theta}_{\text{\tiny Q}}}>\zeta_{\text{\tiny Q}}\}}=1\text{ \ or \ }C_{\text{\tiny M}/\text{\tiny Q}}\leq 1,\\ \end{cases}$ (11)
$\text{with \ }C_{\text{\tiny M}/\text{\tiny Q}}=\left(\frac{T_{\widehat{\theta}_{\text{\tiny Q}}}}{\sqrt{p\log\log(n)}}\right)^{-1}\frac{\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}^{2}}{\log(p\log(n))},$
@@
With large enough constants $\bar{c},c\in(0,\infty)$, set $\zeta_{\text{\tiny M}}=\bar{c}\|\bm{\Sigma}\|^{1/2}\Psi\sqrt{\log(p\log(n))}$ and $\zeta_{\text{\tiny Q}}=c\|\bm{\Sigma}\|\Psi^{2}\sqrt{p\log\log(n)}$, and also $\varpi_{\text{\tiny M}}\asymp\log(p\log(n))$ and $\varpi_{\text{\tiny Q}}\asymp\log\log^{3}(n)$. Suppose that $p\geq n$, and in the case of a change point ($q=1$), we suppose that at least one of (5) and (8) holds. If no change point is present ($q=0$), we regard that $\theta=n$. Then, there exist some constants $\bar{c}_{i}$ and $c_{i}$, $i=0,1,2$ ($\bar{c}_{0}$ and $c_{0}$ appearing in (5) and (8)), which may depend on $\bar{c}$ and $c$, such that,
$\mathsf{P}\left\{\widehat{q}=q\,;\,|\widehat{\theta}_{\text{\tiny O}}-\theta|\leq\|\bm{\Sigma}\|\Psi^{2}\min\left[\bar{c}_{1}|\bm{\Sigma}\bm{\delta}|_{\infty}^{-2}\log(p\log(n)),\,c_{1}|\bm{\Sigma}\bm{\delta}|_{2}^{-2}\sqrt{p\log\log(n)}\right]\right\}\\ \geq 1-\bar{c}_{2}(p\log(n))^{-1}-c_{2}\log^{-1}(n).$
@@
Adapting to unknown sparsity in estimation and inference has been actively explored in the high-dimensional statistics literature. On the change point front, Wang and Feng, (2023) establish the distributions of max- and sum-type statistics for testing for a change in the mean of high-dimensional panel data, as well as their asymptotic independence, which enables sparsity-adaptive testing. For this problem, aggregation based on a cross-sectional grid has also been considered for change point detection (Enikeeva and Harchaoui,, 2019; Liu et al.,, 2021; Pilliat et al.,, 2023; Kovács et al.,, 2024; Moen et al.,, 2024). For linear models, Verzelen and Gassiat, (2018) propose a sparsity-adaptive estimator of the explained variance which achieves optimality by combining a quadratic form of the response and the square-root Lasso estimator.
@@
The change point problem in linear models can be viewed as a special case of the change point problem in covariances, which allows us to place the performance of our proposed methods in a context. This is particularly fruitful in elucidating the role of terms such as $\|\bm{\Sigma}\|$ and $|\bm{\Sigma}^{1/2}\bm{\beta}_{j}|_{2}$ (through $\Psi$) that are often treated as constants in the literature on change point detection in linear models. Compared to the state-of-the-art methods (Wang et al., 2021a, ; Dette et al.,, 2022; Moen,, 2024) that target different types of changes in covariances, our OcScan consistently achieves equal or superior performance across a variety of structural settings. For instance, OcScan attains a sharper detection boundary than the projection-based method of Wang et al., 2021a , with an improvement factor of $p\min(\|\bm{\Sigma}\|\sigma_{*}^{-1}\mathfrak{s},|\bm{\Sigma}\bm{\delta}|_{0},\sqrt{p})^{-1}$. This gain arises from effectively leveraging the specific structure inherent in the linear model. Detailed discussion can be found in Appendix A.

### v3

@@ 3.2 Dense regime
The construction of $T_{k}$ is motivated by the estimator of the explained variance proposed by Dicker (2014) and Verzelen and Gassiat (2018) and indeed, $T_{\theta}$ approximates $\tfrac{\theta(n-\theta)}{n}|\bm{\Sigma}\bm{\delta}|_{2}^{2}$, which follows from the following general result.
@@
In contrast to $\bar{T}_{k}$ in (3), the statistic $T_{k}$ involves the $\ell_{2}$-aggregation of the contrast between the local sample covariance before and after $k$, and thus lends itself naturally to the detection of the change point in the dense regime. We can write $T_{k}=\sum_{i\in[n]}(\gamma_{i}-\bar{\gamma})\bigl(\mathbf{w}_{i}^{\top}\widetilde{\mathbf{Y}}(k)\bigr)^{2}$, where $(\gamma_{i},\mathbf{w}_{i}),\,i\in[n]$, denote the pairs of eigenvalues and eigenvectors of $\mathbf{X}\mathbf{X}^{\top}$ and $\bar{\gamma}=n^{-1}\sum_{i\in[n]}\gamma_{i}$. This representation reveals the importance of centring of $\mathbf{X}\mathbf{X}^{\top}$ in (6), which enables $T_{k}$ to amplify when $\mathsf{E}(\widetilde{\mathbf{Y}}(k)\mid\mathbf{X})$ is more aligned with $\mathbf{w}_{i}$’s associated with larger $\gamma_{i}$’s.
@@
Under the AMOC model (1), let $p\geq n$ and $\varpi_{\text{\tiny Q}}\asymp\bigl(\log\log(n)\bigr)^{3}$. Suppose either that there is a change point, with $q=1$, $\Delta\geq 2\varpi_{\text{\tiny Q}}$ and
@@
$\mathsf{P}\left\{\widehat{q}_{\text{\tiny Q}}=q\,;\,|\widehat{\theta}_{\text{\tiny Q}}-\theta|\leq c_{1}|\bm{\Sigma}\bm{\delta}|_{2}^{-2}\|\bm{\Sigma}\|\Psi^{2}\sqrt{p\log\log(n)}\right\}\geq 1-c_{2}\bigl(\log(n)\bigr)^{-1},$
where $c$ and $c_{i},\,0\leq i\leq 2$, are constants that depend only on $C_{0}$ and $C_{1}$ in Proposition 4.
With $\sigma_{*}$ denoting the smallest non-zero eigenvalue of $\bm{\Sigma}$, it holds that $|\bm{\Sigma}\bm{\delta}|_{2}^{2}\geq\sigma_{*}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}$. Thus, QcScan is minimax optimal in the dense regime where $\mathfrak{s}\geq\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$ (cf. Table 1), provided that $\sigma_{*}^{-1}\|\bm{\Sigma}\|$ is bounded from the above (see also the discussions below Theorem 3). When $\bm{\Sigma}$ is known, pre-rotating $\mathbf{x}_{t}$ allows us to re-write (8) without the dependence on $\bm{\Sigma}$, as
@@
    As in Verzelen and Gassiat (2018), we focus on the high-dimensional setting by assuming that $p\geq n$ in Proposition 4. In the low-dimensional setting with $p<n$, the conclusion of the proposition becomes $\mathsf{P}\{|T_{k}-f_{k}|\geq C_{0}\|\bm{\Sigma}\|\Psi^{2}\sqrt{nz}\}\leq C_{1}e^{-z}$ and accordingly, the results in Theorem 5 no longer depend on $p$. The condition (8) implies that $p\lesssim n^{2}$, which sets the dense regime apart from the sparse regime as in the latter, consistent change point detection is feasible provided that $p=O(\exp(n))$.
@@
    The proof of Proposition 4 is based on the exponential concentration inequalities for Gaussian chaos (Adamczak and Wolff, 2015). We may relax the Gaussianity assumption to permit sub-Weibull distributions by means of Theorems 1.3 and 1.5 of Götze et al. (2021), and thus extend the guarantees of QcScan to non-Gaussian settings.
@@ 3.3.1 When the differential parameter is sparse
Directly imposing sparsity on $\bm{\delta}$, Cho et al. (2025) propose $\ell_{1}$-penalised estimator for its estimation, which is obtained as
@@
Suppose that either (i) $\widehat{\bm{\delta}}$ in (9) is obtained with $\widehat{\theta}=\widehat{\theta}_{\text{\tiny M}}$ and the conditions in Theorem 3 are met, or (ii) $\widehat{\bm{\delta}}$ is obtained with $\widehat{\theta}=\widehat{\theta}_{\text{\tiny Q}}$ and the conditions in Theorem 5 are met. Let $\underline{\sigma}$ denote the smallest eigenvalue of $\bm{\Sigma}$. Provided that $n\geq 4C_{\text{RE}}\underline{\sigma}^{-2}\log(p\vee n)$ and $\lambda=C_{\lambda}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi\sqrt{\log(p\vee n)}$ for some constants $C_{\text{RE}},C_{\lambda}\in(0,\infty)$, we have under (i),
[…]
@@
$\bigl|\widehat{\bm{\delta}}-\bm{\delta}\bigr|_{2}$ $\lesssim\frac{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Psi\max\{\sigma_{X}\sqrt{\mathfrak{s}_{\delta}\log(p\vee n)},\|\bm{\Sigma}\|^{1/2}\bigl(p\log\log(n)\bigr)^{1/4}\}}{\underline{\sigma}\sqrt{\Delta}}\text{ \ and \ }$
$\bigl|\widehat{\bm{\delta}}-\bm{\delta}\bigr|_{1}$ $\lesssim\frac{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\Psi\sqrt{\mathfrak{s}_{\delta}}\max\{\sigma_{X}\sqrt{\mathfrak{s}_{\delta}\log(p\vee n)},\|\bm{\Sigma}\|^{1/2}\bigl(p\log\log(n)\bigr)^{1/4}\}}{\underline{\sigma}\sqrt{\Delta}},$
with probability at least $1-c_{2}\bigl(\log(n)\bigr)^{-1}-c^{\prime}(p\vee n)^{-1}$, for some universal constant $c^{\prime}\in(0,\infty)$ and $\bar{c}_{2}$ and $c_{2}$ in Theorems 3 and 5.
For the consistency of $\widehat{\bm{\delta}}$, we make the stronger assumption requiring $\underline{\sigma}>0$, compared to that required for the consistency in change point detection involving $\sigma_{*}$, the smallest non-zero eigenvalue of $\bm{\Sigma}$. In the inherently sparse regime with $\widehat{\theta}=\widehat{\theta}_{\text{\tiny M}}$, the consistency of $\widehat{\bm{\delta}}$ is analogous to that of the Lasso estimator of the regression parameter (see e.g. Bickel et al., 2009), with $\Delta=\min(\theta,n-\theta)$ taking the role of sample size. Cho et al. (2025) additionally propose a bias-corrected estimator of $\bm{\delta}$ and establish a non-asymptotic Gaussian approximation result, which enables simultaneous inference about all $p$ coordinates of $\bm{\delta}$. Later in Section 4.2, we devise a change point estimator which, based on scanning the projected covariance as $\widehat{\bm{\delta}}^{\top}\mathbf{X}^{\top}\widetilde{\mathbf{Y}}(k)$, attains a refined rate of estimation that is inversely proportional to $\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$.
@@ 3.3.2 For the differential parameter of arbitrary sparsity
Generally, it is infeasible to infer $\bm{\delta}$, or even the regression parameter under stationarity without any assumption on its sparsity. Yet it is still of interest to learn the strength of the change measured by $V_{\circ}:=\bm{\delta}^{\top}\bm{\Sigma}\bm{\delta}$, a quantity closely related to the signal-to-noise ratio or the explained variance in linear models. In particular, $\sigma^{-2}V_{\circ}$ determines the total variation distance between the joint distributions of $\{(Y_{t},\mathbf{x}_{t})\}_{t\in[n]}$ with and without a change, see Lemma 2. Motivated by Proposition 4 and Theorem 5, we propose the following estimator
@@
with probability at least $1-c^{\prime\prime}\bigl(\log(n)\bigr)^{-1}$, with $c^{\prime\prime}\in(0,\infty)$ depending only on $C_{1}$ in Proposition 4.
The upper bound in (10) tends to zero as the gap between the two sides of (8) increases. If $\bm{\Sigma}$ is known, we can pre-rotate the data so that $\widehat{V}$ estimates $V_{\circ}$ directly. In the absence of such knowledge, we conjecture that it is not possible to estimate $V_{\circ}$ without imposing any constraint on the sparsity of $\bm{\delta}$, which is supported by the impossibility result derived in Theorem 4.5 of Verzelen and Gassiat (2018) for the related problem of estimating the explained variance in stationary linear models.
@@ 4.1 Adaptively optimal change point detection
Section 3 demonstrates that McScan and QcScan are minimax (near-)optimal in sparse and dense regimes, respectively. Each regime is determined by whether $\mathfrak{s}<\sqrt{p{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\log\log(n)}}$ or not, where $\mathfrak{s}$ is typically unknown in practice. This calls for a method that adapts to unknown sparsity level. We propose to combine the outputs from McScan and QcScan, which leads to a procedure that detects a change whenever either of the detection requisites in (5) and (8) is met. Further, the resulting change point estimator inherits the localisation property from either $\widehat{\theta}_{\text{\tiny M}}$ or $\widehat{\theta}_{\text{\tiny Q}}$ that attains the better rate of estimation. In short, referred to as OcScan (optimal covariance scanning), the combined method adaptively enjoys the optimality of McScan and QcScan without the explicit knowledge of $\mathfrak{s}$.
It is natural to determine the existence of the change point by $\widehat{q}=\max\{\mathbb{I}_{\{\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}>\zeta_{\text{\tiny M}}\}},\,\mathbb{I}_{\{T_{\widehat{\theta}_{\text{\tiny Q}}}>\zeta_{\text{\tiny Q}}\}}\}$. If $\widehat{q}=0$, we set ${\widehat{\theta}_{\text{\tiny O}}}=n$, while if $\widehat{q}=1$ (with the convention that $\bar{T}_{n}=T_{n}=0$),
[…]
$\text{with \ }C_{\text{\tiny M}/\text{\tiny Q}}=\left(\frac{T_{\widehat{\theta}_{\text{\tiny Q}}}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\|\bm{\Sigma}\|}\sqrt{p\log\log(n)}}\right)^{-1}\frac{\bar{T}_{\widehat{\theta}_{\text{\tiny M}}}^{2}}{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}^{2}}\log(p\log(n))},$
@@
With large enough constants $\bar{c},c\in(0,\infty)$, set $\zeta_{\text{\tiny M}}=\bar{c}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi\sqrt{\log(p\log(n))}$ and $\zeta_{\text{\tiny Q}}=c\|\bm{\Sigma}\|\Psi^{2}\sqrt{p\log\log(n)}$, and also $\varpi_{\text{\tiny M}}\asymp\log(p\log(n))$ and $\varpi_{\text{\tiny Q}}\asymp\bigl(\log\log(n)\bigr)^{3}$. Suppose that $p\geq n$, and in the case of a change point ($q=1$), we suppose that at least one of (5) and (8) holds. If no change point is present ($q=0$), we regard that $\theta=n$. Then, there exist some constants $\bar{c}_{i}$ and $c_{i}$, $i=0,1,2$ ($\bar{c}_{0}$ and $c_{0}$ appearing in (5) and (8)), which may depend on $\bar{c}$ and $c$, such that,
$\mathsf{P}\left\{\widehat{q}=q\,;\,|\widehat{\theta}_{\text{\tiny O}}-\theta|\leq\Psi^{2}\min\left[\bar{c}_{1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}^{2}}|\bm{\Sigma}\bm{\delta}|_{\infty}^{-2}\log(p\log(n)),\,c_{1}\|\bm{\Sigma}\||\bm{\Sigma}\bm{\delta}|_{2}^{-2}\sqrt{p\log\log(n)}\right]\right\}\\ \geq 1-\bar{c}_{2}(p\log(n))^{-1}-c_{2}\bigl(\log(n)\bigr)^{-1}.$
@@
Adapting to unknown sparsity in estimation and inference has been actively explored in the high-dimensional statistics literature. On the change point front, Wang and Feng (2023) establish the distributions of max- and sum-type statistics for testing for a change in the mean of high-dimensional panel data, as well as their asymptotic independence, which enables sparsity-adaptive testing. For this problem, aggregation based on a cross-sectional grid has also been considered for change point detection (Enikeeva and Harchaoui, 2019; Liu et al., 2021; Pilliat et al., 2023; Kovács et al., 2024; Moen et al., 2024). For linear models, Verzelen and Gassiat (2018) propose a sparsity-adaptive estimator of the explained variance which achieves optimality by combining a quadratic form of the response and the square-root Lasso estimator.
@@
The change point problem in linear models can be viewed as a special case of the change point problem in covariances, which allows us to place the performance of our proposed methods in a context. This is particularly fruitful in elucidating the role of quantities such as $\sigma_{X}$, $\|\bm{\Sigma}\|$ and $|\bm{\Sigma}^{1/2}\bm{\beta}_{j}|_{2}$ (through $\Psi$) that are often treated as constants in the literature on change point detection in linear models. Compared to the state-of-the-art methods (Wang et al., 2021a; Dette et al., 2022; Moen, 2024) that target different types of changes in covariances, our OcScan consistently achieves equal or superior performance across a variety of structural settings. For instance, OcScan attains a sharper detection boundary than the projection-based method of Wang et al. (2021a), with an improvement factor of $p\min(\|\bm{\Sigma}\|\sigma_{*}^{-1}\mathfrak{s},|\bm{\Sigma}\bm{\delta}|_{0},\sqrt{p})^{-1}$. This gain arises from effectively leveraging the specific structure inherent in the linear model. In addition, scanning for (univariate) changes in the marginal variance of the response can be beneficial, particularly in regimes with highly unbalanced signal strengths (namely, when $\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{0}\rvert_{2}$ and $\lvert\bm{\Sigma}^{1/2}\bm{\beta}_{1}\rvert_{2}$ differ substantially). Further discussion and comparison can be found in Appendix A.
We finally discuss adaptation to the unknown quantities including $\sigma_{X}^{2}$, $\|\bm{\Sigma}\|$ and $\Psi$. Let us write $\widehat{\bm{\Sigma}}=n^{-1}\sum_{t\in[n]}\mathbf{x}_{t}\mathbf{x}_{t}^{\top}$, from which we derive the approximation of $\sigma_{X}^{2}$ and $\|\bm{\Sigma}\|$ as $\widehat{\sigma}_{X}^{2}=\max_{i\in[p]}\widehat{\Sigma}_{ii}$ and $\|\widehat{\bm{\Sigma}}\|$, respectively, with $\widehat{\Sigma}_{ii},\,i\in[p]$, denoting the diagonal entries of $\widehat{\bm{\Sigma}}$. Also, let $\widehat{\Psi}^{2}_{1,t}=t^{-1}\sum_{u=1}^{t}Y_{u}^{2}$ and $\widehat{\Psi}^{2}_{2,t}=t^{-1}\sum_{u=n-t+1}^{n}Y_{u}^{2}$, and define a dyadic grid $\mathcal{T}=\{2^{\ell}\,:\,\lceil\log_{2}\log\log(n)\rceil\leq\ell\leq\lfloor\log_{2}(n/2)\rfloor\}$, from which we obtain $\widehat{\Psi}^{2}=\max_{t\in\mathcal{T}}(\widehat{\Psi}^{2}_{1,t}\vee\widehat{\Psi}^{2}_{2,t})$. With probability tending to one, we can show that $\widehat{\sigma}_{X}^{2}\asymp\sigma_{X}^{2}$, $\|\widehat{\bm{\Sigma}}\|\asymp\|\bm{\Sigma}\|$ and $\widehat{\Psi}^{2}\asymp\Psi$, leading to:
Corollary 9.
Suppose that $n\leq p\leq c_{*}n$ for some constant $c_{*}\in(0,\infty)$. Then, Proposition 8 continues to hold with $\zeta_{\text{\tiny M}}=\bar{c}\widehat{\sigma}_{X}\widehat{\Psi}\sqrt{\log(p\log(n))}$ and $\zeta_{\text{\tiny Q}}=c\|\widehat{\bm{\Sigma}}\|\widehat{\Psi}^{2}\sqrt{p\log\log(n)}$.
The restrictive condition that $p\leq c_{*}n$ arises from controlling $\|\widehat{\bm{\Sigma}}-\bm{\Sigma}\|$, and it can be lifted when deriving the consistency of McScan with the adaptively selected $\zeta_{\text{\tiny M}}$. Alternatively, we may adopt bootstrap techniques (Chernozhukov et al., 2023) in combination with McScan; a full exploration of such an approach lies beyond the scope of the present paper but represents an interesting direction for future research.